Let T: R^3 to R^3 be the linear transformation defined by


T(x,y,z)= (-x,x-y, 3x+2y+z)


Check whether T satifies the polynomial (x-1)(x+1)^2. Also the find of minimal polynomial of T.



1
Expert's answer
2022-01-12T18:18:21-0500



"T=x\begin{pmatrix}
 -1 \\
 1\\
3
\end{pmatrix}+y\begin{pmatrix}
 0 \\
 -1\\
2
\end{pmatrix}+z\begin{pmatrix}
 0 \\
 0\\
1
\end{pmatrix}"


"T=\begin{pmatrix}
 -1&0 & 0 \\
 1&-1 & 0\\
3&2&1
\end{pmatrix}"




"T-1=\begin{pmatrix}
 -2&0& 0 \\
 1&-2 & 0\\
3&2&0
\end{pmatrix}"



"(T+1)^2=\begin{pmatrix}
 0&0&0 \\
 1&0&0\\
3&2&2
\end{pmatrix}^2=\begin{pmatrix}
 0&0&0\\
 0&0& 0\\
8&4&4
\end{pmatrix}"



"(T-1)(T+1)^2=\begin{pmatrix}
 -2&0& 0 \\
 1&-2 & 0\\
3&2&0
\end{pmatrix}\begin{pmatrix}
 0&0&0 \\
 0&0 & 0\\
8&4&4
\end{pmatrix}=\begin{pmatrix}
 0&0& 0 \\
 0&0& 0\\
0&0&0
\end{pmatrix}"




T satisfies the polynomial "(x-1)(x+1)^2"


"(T-1)(T+1)=\begin{pmatrix}
 -2&0&0 \\
 1&-2 & 0\\
3&2&0
\end{pmatrix}\begin{pmatrix}
 0&0& 0\\
 1&0&0 \\
3&2&2
\end{pmatrix}=\begin{pmatrix}
 0&0 & 0\\
 -2&0&0\\
2&0&0
\end{pmatrix}\ne\begin{pmatrix}
 0&0 & 0 \\
 0&0& 0\\
0&0&0
\end{pmatrix}"




"\therefore" Minimal polynomial is "(x-1)(x+1)^2"